OpenStudy (anonymous):

indefinite integral of 2x+5/cube root of x

OpenStudy (anonymous):

I got the answer when x>0 but i need the general answer

OpenStudy (irishboy123):

that could mean anything, have you tried drawing it ?!?! 🍀

OpenStudy (irishboy123):



OpenStudy (anonymous):

we havent been taught how to draw it

OpenStudy (irishboy123):

scan it!

OpenStudy (anonymous):

the picture is blank

OpenStudy (anonymous):

scan it?

OpenStudy (irishboy123):



OpenStudy (irishboy123):

but in a way that your problem is clear :p

OpenStudy (anonymous):

\[(2x+5)\div \sqrt[3]{x}\]

OpenStudy (anonymous):

the indefinite integral of that

OpenStudy (irishboy123):

you are still using \(\div\) signs?!?! yet doing calculus?!?! that is funny 😀 maybe you mean \(\dfrac{(2x+5)}{ \sqrt[3]{x}}\)??? so break out the exponent rules, ..., and go from there. cheers!

OpenStudy (anonymous):

lets do an example: \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{3x-6}{\sqrt[5]{x}}{\rm~d}x}\) \(\large\color{black}{\displaystyle\sqrt[5]{x}=\sqrt[5]{x^1}=x^{1/5}}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{3x-6}{x^{1/5}}{\rm~d}x}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{3x}{x^{1/5}} - \frac{6}{x^{1/5}}{\rm~d}x}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{3x}{x^{1/5}}{\rm~d}x - \int\limits_{~}^{~}\frac{6}{x^{1/5}}{\rm~d}x}\) \(\large\color{black}{\displaystyle 3x \div x^{1/5}=3x^1 \div x^{1/5}=3x^{1~-~1/5}=3x^{4/5}}\) \(\large\color{black}{\displaystyle 6 \div x^{1/5}=6 \times x^{-1/5}=6 x^{-1/5}}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}3x^{4/5}{\rm~d}x - \int\limits_{~}^{~}6 x^{-1/5}{\rm~d}x}\) \(\large\color{green}{\displaystyle \frac{1}{4/5~+1}\cdot 3\cdot x^{4/5~+1} {~~-~~} \frac{1}{-1/5~+1}\cdot6 \cdot x^{-1/5} }\) \(\large\color{green}{\displaystyle \frac{1}{9/5}\cdot 3\cdot x^{9/5} {~~-~~} \frac{1}{4/5}\cdot6 \cdot x^{4/5}}\) \(\large\color{green}{\displaystyle \frac{5}{9}\cdot 3\cdot x^{9/5} {~~-~~} \frac{5}{4}\cdot6 \cdot x^{4/5} }\) \(\large\color{green}{\displaystyle \frac{5}{3} x^{9/5} {~~-~~} \frac{15}{2} x^{4/5} }\) \(\large\color{red}{\displaystyle \frac{5}{3} x^{9/5} - \frac{15}{2} x^{4/5}+C }\)